6. Use given data to calculate the standard free- energy change for the reaction Pa(g) + 6 Cl2(g) → 4 PC13(g) at 298 K. (ii)Need help with #6,7&9 to make sure I did it right. Will rate high if it is correct! Thank you.
Answer #1

tu (g) + 6C ca 4 Paz eg) A4 = 544) Product - Al Roactant = 4x (-299-6) - 610) - 94.4 = - 1222-8 kJ / moles 19%eovnse = -05 1